OpenStudy (anonymous):

Which of the following are point-slope equations of the line going through (-5, 6) and (1, 1)? Check all that apply, because there's more than one answer . A.y - 1 = -6/5 (x - 1) B.y - 6 = -6/5 (x + 5) C.y + 1 = 5/6 (x - 1) D.y + 6 = -6/5 (x - 5) E.y - 6 = -5/6 (x + 5) F.y - 1 = -5/6 (x - 1)

OpenStudy (whpalmer4):

First thing you need to do is find the slope of a line through those points. \[m=\frac{y_2-y_1}{x_2-x_1}\]will give you the slope. After you have the slope, remember that the point-slope equation for a line through a known point \((x_0,y_0)\) is \[y-y_0 = m(x-x_0)\] Armed with that, you should be able to quickly decide which equations meet the constraints.

OpenStudy (whpalmer4):

Note that you want to write out an equation for each of the two points you know...

Directrix (directrix):

m = (6-1)/(-5-1) = ?

OpenStudy (anonymous):

0.06 @Directrix

Directrix (directrix):

m = 5/(-6) Check and see.